OpenStudy (anonymous):

Solve triangle ABC. (If an answer does not exist, enter DNE. Round your answers to one decimal place.) PLEASE HELP!! a = 67, c = 50, ∠C = 54° ∠A = ° ∠B = ° b =

OpenStudy (anonymous):

What is your question? What are you looking for?

OpenStudy (anonymous):

It is basally asking to solve for the triangle with the given information. with the angles make a triangle out of it and find the angles..but i need help with that

OpenStudy (anonymous):

Well the sum of all internal angles is 180 degrees.For B lets use the formula B=180-(A+C)

OpenStudy (anonymous):

After you solve that use the sine rule.b/sinB and so on

OpenStudy (anonymous):

Do you know the Law of Sine?

OpenStudy (anonymous):

isn't sin opposite over hypotnuse?

OpenStudy (anonymous):

Well that's sine but not the Law of sine

OpenStudy (anonymous):

\[\frac{ a }{ \sin(A) }=\frac{ b }{ \sin(B) }=\frac{ c }{ \sin(C) }\] Have you learned that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

 not draw to scale

OpenStudy (anonymous):

Ok so, we can use the law of Sine to solve for angle A Using the formula: \[\frac{ 50 }{ \sin(54) } = \frac{ 67 }{ \sin(C) }\] Now you have to solve for sin(C), do you know how to do this?

OpenStudy (anonymous):

its the same way as shown above